Suppose that the price of commodity Y is $ 20 per unit while the price of commodity X is $ 15 

per unit and suppose that an individual’s money income is $ 100 per period and is all spent on X & Y. 

i. Draw the budget constraint line for this consumer at the initial point. ​​


1
Expert's answer
2021-05-30T15:32:25-0400

For the graphical representation of the budget line, represent good Y on the vertical axis and the good X on the horizontal axis.

The intercept of vertical axis is calculated as follows:

intercept on the vertical axis"=\frac{income}{price \space of\space good \space Y}"

"=\frac{100}{20}=5"


The intercept of horizontal axis is calculated as follows:

intercept on the horizontal axis"=\frac{income}{price \space of\space good \space X}"

"=\frac{100}{15}=6.67"

Joining the vertical and horizontal intercept we get budget constraint as shown in figure below:


Budget line is linear downward sloping.